# Azumaya algebras and Barr’s Theorem

### v3

@@ Introduction
The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of strictly Henselian local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following [^1] We study these characterisations in a constructive setting in the reference [CLN].:
@@
    an algebra $A$ over $F$ which is finite and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{F}(A)$ is an isomorphism; or, equivalently,
@@
This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$. The definition is a generalisation of the first characterisation: an algebra $A$ which is projective of finite type over $R$ and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{R}(A)$ is an isomorphism. In this note, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
@@ 1 Some coherent theories
We will consider some coherent theories [Joh] in the language of rings.
@@ 1.1 Theory of local rings
$0\neq 1~~~~~~~~~\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
@@
If we add to our language a new predicate $\mathsf{m}(x)$ with the axioms
$\mathsf{inv}(x)\vee\mathsf{m}(x)~~~~~~~~~~\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
then we get the theory of local rings that are residually discrete [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper., i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
It does not seem possible however to express the property of being Henselian in a coherent way in the language of rings (i.e. without introducing the predicate $\mathsf{m}(x)$). Surprisingly, as noticed by G. Wraith [Wraith], the property of being strictly Henselian, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way. In this note, we present some remarks connected to this observation. The first one is a small variation of Wraith’s axiomatisation. Both axiomatisations can be used to define the étale site over a ring. Finally, we show in a constructive setting that Azumaya algebras over a ring are exactly the algebras that are locally matrix algebras for the étale topology.
@@ 1.2 Theory of separably closed local rings
We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D:R\rightarrow Z(R)$ over $R$. A support is a map $d:R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
$d(1)=1~~~~~d(0)=0~~~~~~d(a+b)\leqslant d(a)\vee d(b)~~~~~~~~d(ab)=d(a)\wedge d(b)$
We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\dots\vee d(a_{n})$. It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$. Conversely, it is elementary to check that we get a support if we define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$, and this is a concrete way to build the universal support.
@@
$d(a_{1},\dots,a_{n})=d(s_{1},\dots,s_{n})\text{,}$
where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
For the formulation of this theory of strictly Henselian rings, we need the notion of universal splitting algebra [LQ, III, 4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\dots$
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~\dots,~u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know, using Cauchy modules [LQ, III, 4.3], that $L=R[x_{1},\dots,x_{n}]$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
For the notion of faithfully flat extension, we follow the definition in [LQ, VIII, 6]: it is an $R$-algebra $i:R\rightarrow S$ such that
  • if $u$ is a row vector in $A^{n}$ and $v$ a column vector in $B^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $A$ and $w$ such that $uA=0$ and $v=i(A)w$;
  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $B$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $A$.
It then follows [LQ, VIII, 6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finite free $R$-module of rank >0>0.
@@
We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$. Since $D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=D(\delta_{1}(f),\dots,\delta_{n}(f))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in $L$.
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$. This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $F$.
@@
Let $R$ be a local ring, residually discrete with maximal ideal $\mathsf{m}$ and residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra AA, then $f$ is unramifiable.
Lemma 1.2.
If $f=(X-a_{1})\dots(X-a_{p})g$ in $R[X]$ with $g$ monic of degree $n$ and $\Delta(f)=1$ and $R$ is local, then $f^{\prime}(a_{1})$ is invertible or $\dots$ or $f^{\prime}(a_{p})$ is invertible or $\Delta(g)=1$.
@@
We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $g$,
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
@@
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),g^{\prime}(x_{1}),\dots,g^{\prime}(x_{n}))$
@@
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
@@
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
@@
Lemma 1.3.
If $R$ is local and any monic unramifiable polynomial has a root in $R$ then any monic unramifiable polynomial has a simple root in $R$.
@@
Assume that $R$ is local and any monic unramifiable polynomial has a root in $R$. Let $f$ be a monic unramifiable polynomial. It has a root $a_{1}$ and we can write $f=(X-a_{1})f_{1}$. By Lemma 1.2, $a_{1}$ is a simple root of $f$ or $\Delta(f_{1})=1$. If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $f=(X-a_{1})(X-a_{2})f_{2}$. Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $f$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $f$. ∎
Let $k$ be a ring. We consider now the category of finitely presented $k$-algebras with the covering
@@
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
It also follows from Lemma 1.3 that we can instead take the following covering, since it classifies an equivalent theory.
@@
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(P)[1/P^{\prime}]$ if $P$ is monic and unramifiable.
This defines a site, in the sense of Grothendieck (see e.g. [Joh]), over the opposite of the category of finitely presented $k$-algebras. The corresponding sheaf model is the classifying topos of the theory of strictly Henselian $k$-algebras [Joh]. In this sheaf model, we have a “generic” $k$-algebra which is strictly Henselian.
@@
Theorem 1.1.
If $R$ is a local ring which is also residually discrete, then $R$ is strictly Henselian if, and only if, $R$ is Henselian and its residue field is separably closed.
@@
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write P=(Xa)QP=(X-a)Q, and $u$ is residually a simple root of $Q$. ∎
@@ 2 Some lemmas about Henselian rings
In the next section, we need some results about Henselian rings. In order to make the paper self-contained we list here these results with their proofs. They are taken from the reference [ALP], with a variation in the proof of Lemma 2.1.
We assume in this section that $R$ is local residually discrete with maximal ideal $\mathsf{m}$ and Henselian.
Lemma 2.1 ([ALP, Prop. 5.1]).
Let $P$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(P)$. If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod. $\mathsf{m}R[x]$, then we can find $u$ idempotent in $R[x]$ and $u=e(x)$ mod. $\mathsf{m}R[x]$.
@@
Let $A=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 0 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod. $\mathsf{m}$. We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$. We let [^3] The $e_{I},~|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod. $\mathsf{m}$. $k<n$ be such that $e_{I}\neq 0$ mod. $\mathsf{m}$ if $|I|=k$ and $e_{I}=0$ mod. $\mathsf{m}$ if $|I|=k+1$. Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod. $\mathsf{m}$. We have $\Sigma_{|I|=k}e_{I}=1$ mod. $\mathsf{m}$ since it is an element in $R$, which is idempotent mod. $\mathsf{m}$. Hence the family of the $e_{I},~|I|=k$, forms an FSOI [^4] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sec. II-4]). mod. $\mathsf{m}A$.
@@
$Q(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathsf{m}R[X]$ with $N=\binom{n}{k}$.
Since $R$ is Henselian, we have $\alpha$ in $R$ such that $Q(\alpha)=0$ and $\lambda=Q^{\prime}(\alpha)$ is invertible. If we write $Q_{I}=Q/(X-e_{I})$ then $Q^{\prime}(X)=\Sigma_{I}Q_{I}$. We take
$v_{I}=\frac{1}{\lambda}Q_{I}(\alpha)$
and we then have that the $v_{I},~|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$. We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod. $\mathsf{m}A$. Let $L_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$. If we write $u_{1}$ for $\Sigma_{I\in L_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in L_{1}}e_{I}=e(x_{1})$ mod. $\mathsf{m}A$. This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in L_{1}}e_{I}$ mod. $\mathsf{m}A$. Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathsf{m}A=\mathsf{m}R[x_{1}]$. ∎
For completeness, we recall the statement and proof of Nakayama’s Lemma [LQ, Lemma IX-2.1].
Lemma 2.2.
Let $M$ be a finitely generated $R$-module and $N$ an $R$-submodule such that $M=N+\mathsf{m}M$, then $M=N$.

### v4

# arXiv:2306.17679v4[math.AC] 01 Mar 2026

@@ Introduction
The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of separably closed local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
A fundamental invariant associated to a commutative field $F$ is its Brauer group, whose elements are the division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following: [^1] We study these characterisations in a constructive setting in the reference [CLN].
@@
    an algebra $A$ over $F$ which is finitely generated and such that the canonical map $A\otimes_{F}A^{\mathrm{op}}\rightarrow\operatorname{End}_{F}(A)$ is an isomorphism; or, equivalently,
@@
This notion of central simple algebra over a field has been generalised to the notion of Azumaya algebra over an arbitrary commutative ring $k$. The definition is a generalisation of the first characterisation: an algebra $A$ which is finitely generated projective over the ring $k$ and such that the canonical map $A\otimes_{k}A^{\mathrm{op}}\rightarrow\operatorname{End}_{k}(A)$ is an isomorphism. In this paper, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
@@ 1 Some coherent theories
We consider some coherent theories [Joh] in the language of rings.
@@ 1.1 Theory of local rings
$0\neq 1\qquad\qquad\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
@@
If we add to our language a new predicate $\mathfrak{m}(x)$ with axioms
$\mathsf{inv}(x)\vee\mathfrak{m}(x)\qquad\qquad\neg(\mathfrak{m}(x)\wedge\mathsf{inv}(x))$
then we get the theory of local rings that are residually discrete, [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper. i.e. where the property of being invertible is now decidable, and the predicate $\mathfrak{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $f$, if we have $u$ such that $f(u)\in\mathfrak{m}$ and $f^{\prime}(u)\notin\mathfrak{m}$ then we can find $\alpha$ such that $f(\alpha)=0$ and $\alpha-u\in\mathfrak{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathfrak{m}$ has a root which is equal to $1$ mod. $\mathfrak{m}$. The work [ALP] shows in a constructive setting for this theory the lifting property for idempotents of a finitely generated commutative algebra over a Henselian local ring.
It does not seem possible however to express the property of being Henselian in a coherent way in the language of local rings (i.e. without introducing the predicate $\mathfrak{m}(x)$). Surprisingly, as observed by G. Wraith [Wraith], the property of being separably closed, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way. In this paper, we present some remarks connected to this observation. The first one is a small variation of Wraith’s axiomatisation. Both axiomatisations can be used to define the étale site over a ring. Finally, we show in a constructive setting that Azumaya algebras are exactly the algebras that are locally matrix algebras for the étale topology.
@@ 1.2 Theory of separably closed local rings
We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D\colon R\rightarrow Z(R)$ over $R$. A support is a map $d\colon R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
$d(1)=1\qquad\qquad d(0)=0\qquad\qquad d(a+b)\leqslant d(a)\vee d(b)\qquad\qquad d(ab)=d(a)\wedge d(b)$
We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\cdots\vee d(a_{n})$. It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$. Conversely, it is elementary to check that we get a support if we define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$, and this is a concrete way to build the universal support.
@@
$d(a_{1},\dots,a_{n})=d(c_{1},\dots,c_{n})\text{,}$
where $c_{1},\dots,c_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
For the formulation of the theory of separably closed local rings, we need the notion of universal splitting algebra [LQ, Sect. III-4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\cdots$
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1}$, …, $u_{n}-\sigma_{n}$, with $\sigma_{1}$, …, $\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ the elementary symmetric polynomials. We know, using Cauchy modules [LQ, Fact III-4.3], that $L$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
For the notion of faithfully flat extension, we follow the definition in [LQ, Sect. VIII-6]: it is a commutative $R$-algebra $i\colon R\rightarrow S$ such that
  • if $u$ is a row vector in $R^{n}$ and $v$ a column vector in $S^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $R$ and $w$ such that $uM=0$ and $v=i(M)w$;
  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $S$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $R$.
It then follows [LQ, Sect. VIII-6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finitely generated free $R$-module of rank >0>0.
@@
We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$. Since $\Delta(f)=D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in the universal splitting algebra $L$.
Over a residually discrete local ring with maximal ideal $\mathfrak{m}$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of the residue field $R/\mathfrak{m}$. This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $R/\mathfrak{m}$.
@@
Let $R$ be a local ring, residually discrete with maximal ideal $\mathfrak{m}$ and residue field $R/\mathfrak{m}$. If $g$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial commutative $R/\mathfrak{m}$-algebra, then $g$ is unramifiable.
Lemma 1.2 ([LQ2, Lemme IX-7.1]).
If $g=(X-a_{1})\cdots(X-a_{p})f$ in $R[X]$ with $f$ monic of degree $n$, $\Delta(g)=1$, and $R$ is local, then $g^{\prime}(a_{1})$ is invertible or … or $g^{\prime}(a_{p})$ is invertible or $\Delta(f)=1$.
@@
We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $f$,
$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}),h(x_{1})f^{\prime}(x_{1}),\dots,h(x_{n})f^{\prime}(x_{n}))$
@@
$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}),f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$
@@
$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}))\vee\Delta(f)$
@@
$1=D(g^{\prime}(a_{1}),\dots,g^{\prime}(a_{p}))\vee\Delta(f)$
@@
Definition 1.1 ([LQ2, Définition IX-7.2]).
A local ring $R$ is separably closed if every unramifiable monic polynomial has a root in $R$.
Lemma 1.3 ([LQ2, Lemme IX-7.3]).
If $R$ is a separably closed local ring then any unramifiable monic polynomial has a simple root in $R$.
@@
Assume that $R$ is separably closed local. Let $g$ be a monic unramifiable polynomial. It has a root $a_{1}$ and we can write $g=(X-a_{1})f_{1}$. By Lemma 1.2, $a_{1}$ is a simple root of $g$ or $\Delta(f_{1})=1$. If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $g=(X-a_{1})(X-a_{2})f_{2}$. Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $g$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $g$. ∎
Let $k$ be a ring. We can consider the following two coherent theories for a commutative $k$-algebra $R$:
@@
    $T_{1}$ states that $R$ is local, and that any monic unramifiable polynomial has a root.
@@
    $T_{2}$ states that $R$ is local, and that any monic unramifiable polynomial has a simple root.
We have just seen that these two theories are equivalent.
The equivalence between these two theories can be seen as a logical formulation of the “trick” attributed to O. Gabber described in [Stacks, Tag 02LI].
A morphism
$S\longrightarrow S[X]/(f),$
with $f$ monic, is finite locally free, and thus represents the “finite locally free” stage in Gabber’s refinement of an étale covering. Étaleness is obtained by further localising, that is, by passing to the Zariski covering
$S\longrightarrow S[X]/(f)[1/f^{\prime}].$
Lemma 1.3 provides the logical ingredient of this refinement: over a local base ring, the existence of a root of a monic unramifiable polynomial implies the existence of a simple root. Internally, this shows that coverings generated by finite locally free morphisms of the above form can be refined, after a Zariski localisation, to étale morphisms. In this sense, the equivalence between the two theories $T_{1}$ and $T_{2}$ is the internal counterpart of Gabber’s trick: finite locally free coverings suffice to generate the same local objects as étale coverings, up to Zariski refinement.
These two coherent theories define two different sites (notions of covering), in the sense of Grothendieck (see e.g. [Joh]), over the opposite of the category of finitely presented commutative $k$-algebras [Joh]. The first theory $T_{1}$ corresponds to:
@@
    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
@@
    $S$ is covered by $S\rightarrow S[X]/(f)$ if $f$ is monic and unramifiable.
The second theory $T_{2}$ corresponds to:
  1. 1.
    $S$ is covered by $\rightarrow S[1/s_{1}], …, SS$$\rightarrow S[1/s_{n}$] if $1=(s_{1},\dots,s_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(f)[1/f^{\prime}]$ if $f$ is monic and unramifiable.
Both define the same corresponding classifying topos of the theory of separably closed local commutative $k$-algebras [Joh]. In this sheaf model, we have a “generic” local commutative $k$-algebra $R$ which is separably closed.
@@
Theorem 1.1 ([LQ2, Lemmes IX-7.4, IX-7.5]).
If $R$ is a local ring which is also residually discrete, then $R$ is separably closed if, and only if, $R$ is Henselian and its residue field is separably closed.
@@
Let $g$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $g$ that $g$ has a root in $R$ which is equal to $u$ mod. $\mathfrak{m}$. The polynomial $g$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathfrak{m}$ then we write g=(Xa)fg=(X-a)f, and $u$ is residually a simple root of $f$. ∎
Thus a residually discrete local ring is separably closed if and only if it is “strictly Henselian”.
@@ 2 Some lemmas about Henselian rings
In the next section, we need a generalisation of a result about Henselian rings. In order to make the paper self-contained we mention here this result, Theorem 2.1, taken from the reference [ALP].
We assume in this section that $R$ is local residually discrete with maximal ideal $\mathfrak{m}$ and Henselian.
Lemma 2.1 is a main step in the proof of Theorem 2.1: we provide a variation of the proof that requires less constructive Galois theory than the original one in [ALP].
Lemma 2.1 ([ALP, Prop. 5.7]).
Let $f$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(f)$. If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod. $\mathfrak{m}R[x]$, then we can find $u$ idempotent in $R[x]$ with $u=e(x)$ mod. $\mathfrak{m}R[x]$.
@@
Let $L=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $f$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathfrak{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathfrak{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathfrak{m}$, so it is $1$ or 0 mod. $\mathfrak{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathfrak{m}R[x_{1}]$ and we can take $u_{1}=1$.
We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod. $\mathfrak{m}$. We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$. We let [^3] The $e_{I},~|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod. $\mathfrak{m}L$. $k<n$ be such that $e_{I}\neq 0$ mod. $\mathfrak{m}L$ if $|I|=k$ and $e_{I}=0$ mod. $\mathfrak{m}L$ if $|I|=k+1$. Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod. $\mathfrak{m}L$. We have $\Sigma_{|I|=k}e_{I}=1$ mod. $\mathfrak{m}$ since it is an element in $R$ which is idempotent mod. $\mathfrak{m}$. Hence the family of the $e_{I},~|I|=k$, forms an FSOI [^4] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sect. II-4]). mod. $\mathfrak{m}L$.
@@
$g(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathfrak{m}R[X]$ with $N=\binom{n}{k}$.
Since $R$ is Henselian, we have $\alpha$ in $R$, $\alpha=1$ mod. $\mathfrak{m}$, such that $g(\alpha)=0$; furthermore $\lambda=g^{\prime}(\alpha)$ is invertible with $\lambda=1$ mod. $\mathfrak{m}$. If we write $g_{I}=g/(X-e_{I})$ then $g^{\prime}(X)=\Sigma_{I}g_{I}$. We take
$v_{I}=\frac{1}{\lambda}g_{I}(\alpha)$
and we then have that the $v_{I},~|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$. We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod. $\mathfrak{m}L$. Let $\mathcal{I}_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$. If we write $u_{1}$ for $\Sigma_{I\in\mathcal{I}_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in\mathcal{I}_{1}}e_{I}=e(x_{1})$ mod. $\mathfrak{m}L$. This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in\mathcal{I}_{1}}e_{I}$ mod. $\mathfrak{m}L$. Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathfrak{m}L=\mathfrak{m}R[x_{1}]$. ∎
Theorem 2.1 ([ALP, Theorem 5.9]).
If $A$ is a commutative $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathfrak{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathfrak{m}A$.
The following generalisation is not in [ALP] and will be crucial in the proof of Lemma 3.1.
Corollary 2.1.
If $A$ is a (not necessarily commutative) $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathfrak{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathfrak{m}A$.